# Efficient Retrieval with Learned Similarities

### v1

@@ 1 Introduction
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022). In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
@@
Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020). As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups (Tay et al., 2022; Wang et al., 2022). Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021). Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
@@
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018; Zhai et al., 2023), while large scale recommendation systems commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019) with low end-to-end latency. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023). What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface for these techniques makes it even more difficult for designing a general solution for efficient retrieval.
@@
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022), and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Gao et al., 2021; Zhai et al., 2024).
@@
This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL. To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities. We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions. We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound. Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014). Importantly, our approach efficiently utilizes modern accelerators like GPUs due to higher arithmetic intensity. We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
@@ 2 Expressiveness of Mixture of Logits
$P$ ($P_{q}$, $P_{x}$) number of component-level embeddings. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of component-level embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
@@
MoL applies an adaptive gating weight $\pi_{p}(q,x)$ to $P$ groups of component-level dot products $\langle f_{p}(q),g_{p}(x)\rangle$. Following Zhai et al. (2023):
@@
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings (where $P_{q}\times P_{x}=P$) and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
@@
We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper as the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
@@
Now we show that any high rank matrix can be decomposed into a mixture of logits with low rank matrices. Specifically, without loss of generality, we prove the following:
@@
First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$:
@@
Let $\lambda_{min}=\min{\lambda_{ij}}$, and $\lambda_{max}=\max{\lambda_{ij}}$. Let $B_{1}=\lambda_{min}UI_{d}V$, $B_{2}=\lambda_{max}UI_{d}V$, where $I_{d}$ is the identity matrix of rank $d$. We have $A^{\prime}_{ij}=\lambda_{ij}\sum_{k=1}^{d}U_{ik}V_{kj}=\pi(i,j)\cdot B_{1ij}+(1-\pi(i,j))\cdot B_{2ij}$, where $\pi(i,j)=\frac{\lambda_{max}-\lambda_{ij}}{\lambda_{max}-\lambda_{min}}$. Because $\lambda_{min}\leq\lambda_{ij}\leq\lambda_{max}$, we have $0\leq\pi(i,j)\leq 1$.
@@ 3.1 Exact algorithm
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL score of the items in the set $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$ exhaustively, and return the top $K$ items with the highest scores (line 12).
@@ 3.2 Approximate algorithms
In both heuristics, we perform a two-stage retrieval as shown in Algorithm 2. In the first stage, we retrieve a set of $K^{\prime}$ candidate items that are potentially high in MoL score by using dot products (line 2). Note that $K^{\prime}$ can be larger than $K$, e.g., due to oversampling. In the second stage, we evaluate the MoL score of the candidate items and return the top $K$ items (line 3).
@@ 4 Evaluation
In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3. Our code and pretrained model checkpoints are publicly available at https://github.com/bailuding/rails.
@@ 4.1 Setup
We consider a classical retrieval task (next-item prediction) in recommendation systems (Covington et al., 2016; Zhu et al., 2018; Li et al., 2019; Zhai et al., 2023). We employ three widely used datasets, MovieLens (Harper and Konstan, 2015) (the 1M and 20M subsets), and Amazon Reviews (McAuley et al., 2015) (the largest Books subset). We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets (Kang and McAuley, 2018; Zhai et al., 2023). In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step. Table 2 shows the statistics of these three workloads.
Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ component-level embedding dimension
@@
Model training.
We implemented MoL on top of two popular self-attention sequential encoder architectures, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024). The models are implemented in PyTorch and are trained with 2x NVIDIA RTX 6000 Ada GPUs.
@@
ML-1M SASRec .0610 .2806 .5437 .7498 .1307
SASRec + MoL .0672 .3019 .5711 .7650 .1431
HSTU .0774 .3306 .5897 .7832 .1571
HSTU + MoL .0875 .3412 .6013 .7877 .1692
ML-20M SASRec .0674 .2889 .5503 .7661 .1388
SASRec + MoL .0731 .3068 .5645 .7768 .1498
HSTU .0965 .3556 .6143 .8074 .1800
HSTU + MoL .1014 .3661 .6234 .8116 .1876
@@
We use Recall (Hit Rate) as the main metric. It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3). Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 21.4% in HR@1, an average of 13.7% in HR@10, and an average of 16.2% in MRR across the six scenarios, validating the importance of learned similarity functions. We evaluate our proposed retrieval algorithms relative to the best performing rows, specifically “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the following section.
@@ 4.2 Top $K$ retrieval performance
Table 4 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023). We set batch size to $32$ for ML-1M, ML-20M, and Books. We perform evaluation on a single NVIDIA RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.

### v2

@@ 1 Introduction
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items [5, 9, 26, 25, 2]. In many concrete use cases, such as vector databases [19], the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
@@
Despite the popularity of dot products and numerous work done to improve their efficiency [18, 35, 28, 4], state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values [26, 22]. As another example, Probabilistic Label Trees (PLTs) [16] and Tree-based Deep Models (TDMs) [45, 46] map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups [38, 40]. Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products [8]. Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms [14, 41, 34, 42].
@@
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps [45, 42], while recommendation and NLP deployments commonly need to handle billions of items [6, 26, 3] with low end-to-end latency. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force [14] or heuristics [42]. What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult for designing a general solution for efficient retrieval.
@@
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases [19, 4], and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data [5, 26, 8, 43].
@@
This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL. To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities. We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions. Compared with prior work on learned similarities [14, 38] which often fail to outperform dot product baselines in large-scale settings [33, 36], our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on corpus with hundreds of millions to billions of items [42, 3]. We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound. Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS [18, 35, 19, 12]. We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms. Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity [42], which results in MIPS-level inference latency and throughput. This concludes both theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
@@ 2 Expressiveness of Mixture of Logits
$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ low-rank embeddings ("component-level embeddings") to represent $q$ and $x$. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
@@
$d_{P}$ dimensionality of low-rank (component-level) embeddings. $f_{p}(q),g_{p}(q)\in\mathbb{R}^{d_{P}}$.
@@
MoL [42, 3] assumes that the query $q$ and the item $x$ are already mapped to $P$ groups of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively. MoL then calculates the similarity between the query $q$ and the item $x$ by applying adaptive gating weights, $\pi_{p}(q,x)\in[0,1]$, to the inner products of these $P$ pairs of low-rank embeddings, or $\langle f_{p}(q),g_{p}(x)\rangle$s. Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ [42, 3], and parameterizes $\pi_{p}(q,x)$ utilizing some neural network with $f_{p}(q)$s, $g_{p}(x)$s, and $\langle f_{p}(q),g_{p}(x)\rangle$s as input features, but these details do not affect our analyses in this paper. Following [42]:
@@
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embedding, where $P_{q}\times P_{x}=P$, and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
@@
We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper, given that the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
@@
Now we show that any high rank matrix can be decomposed into a mixture of logits based on low rank matrices. Specifically, without loss of generality, we prove the following:
@@
We can think about $n$ as the number of queries and $m$ the number of items (or vice versa). First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$ ($d\leq n$):
@@
Let $\lambda_{min}=\min{\lambda_{ij}}$, and $\lambda_{max}=\max{\lambda_{ij}}$. Let $B_{1}=\lambda_{min}U{D_{n,d}}V$, $B_{2}=\lambda_{max}U{D_{n,d}}V$, where $D_{n,d}$ denotes an $n$-by-$n$ diagonal matrix with the first $d$ elements of the diagonal being $1$s and the rest being 00s. We have $A^{\prime}_{ij}=\lambda_{ij}\sum_{k=1}^{d}U_{ik}V_{kj}=\pi(i,j)\cdot B_{1ij}+(1-\pi(i,j))\cdot B_{2ij}$, where $\pi(i,j)=\frac{\lambda_{max}-\lambda_{ij}}{\lambda_{max}-\lambda_{min}}$. Because $\lambda_{min}\leq\lambda_{ij}\leq\lambda_{max}$, we have $0\leq\pi(i,j)\leq 1$.
@@ 3.1 Exact algorithm
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL scores of the items in $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$, and return the top $K$ items with the highest scores (line 12).
@@ 3.2 Approximate algorithms
In both heuristics, we perform a two-stage retrieval as shown in Algorithm 2. In the first stage, we retrieve a set of $K^{\prime}$ candidate items that are potentially high in MoL score by using dot products (line 2). Note that $K^{\prime}$ can be larger than $K$, e.g., due to oversampling. In the second stage, we evaluate the MoL scores of the candidate items and return the top $K$ items (line 3).
@@ 4 Evaluation
In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3.
@@ 4.1 Setup
We consider a classical retrieval task (next-item prediction) in recommendation systems [5, 45, 26, 42]. We employ three widely used datasets, the 1M and 20M subsets of MovieLens [13], and the largest Books subset of Amazon Reviews [29]. We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets [20, 42]. In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step, which is then mapped to $P_{q}$ embeddings using an multi-layer perceptron (MLP) when using the MoL learned similarity, consistent with prior work. We discuss detailed hyperparameter settings in Appendix A. Table 2 shows the statistics of these three workloads.
Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ $d_{P}$
@@
Model.
We implemented MoL on top of two popular sequential architectures, SASRec [20] and HSTU [43], with cosine similarity as the default similarity function (SASRec / HSTU rows in Table 3). The models are are implemented in PyTorch and trained with 2x NVIDIA RTX 6000 Ada GPUs.
@@
ML-1M SASRec .0610 .2818 .5470 .7540 .1352
SASRec + MoL .0710 .3034 .5655 .7680 .1448
HSTU .0750 .3332 .5956 .7824 .1579
HSTU + MoL .0837 .3460 .6048 .7874 .1673
ML-20M SASRec .0653 .2883 .5484 .7658 .1375
SASRec + MoL .0790 .3124 .5685 .7784 .1551
HSTU .0962 .3557 .6146 .8080 .1800
HSTU + MoL .1008 .3697 .6244 .8124 .1880
@@
We use Recall (Hit Rate) as the main metric. It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3). Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 24.2% in HR@1, 14.3% in HR@10, and 16.3% in MRR across the six scenarios, validating that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable. We evaluate our proposed retrieval algorithms relative to the best performing rows, “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the next section.
@@ 4.2 Top $K$ retrieval performance
Table 4 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms [19, 4, 42]. We set batch size to $32$ for ML-1M, ML-20M, and Books. We perform evaluation on a single NVIDIA RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.
@@
5 Related work
Similarity Functions in Retrieval.
Most information retrieval models in recommendations and natural language processing follow a classical two-stage paradigm [5], where up to billions of items [6, 26, 3] are first filtered down to hundreds in the retrieval stage, followed by the ranking stage that produces the final top 1-10 candidates. Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function [5, 21]. Researchers quickly realized that dot products limited retrieval stage’s performance, and explored various learned similarity-based approaches. Prominent variants include maximum similarity based on multiple embeddings [26, 22, 34], specialized neural networks, often based on Hadamard products [14, 37, 39, 3], and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal [16, 45, 46] or explicitly in the “generative retrieval” setups [8, 38, 40, 36]. It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendations [33] and DSI in NLP [36]. Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items [3]. Despite these challenges, significant gains (17% gains at Hit Rate@100 [42] to 24% gains at Hit Rate@400 [3]) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions [34, 42], implicit diversification done as part of beam search [8, 36], explicit representation of side-information using special neural architectures [3], and hardware-aware similarity function and inference algorithm design on GPUs [4, 42, 3].
Efficient Nearest Neighbor Search (NNS).
Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures [1, 32], locality sensitive hashing [10, 35], production quantization [18, 11], data partitioning [27, 44], graph-based methods [28, 17], and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results [46]. Algorithms based on implicit [28, 17, 39, 31] or explicit graphs [39] have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case [30]. Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques [35, 11] no longer fully utilize GPUs; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work [4, 42, 31, 3].
6 Conclusion
We have analyzed techniques for efficient retrieval with learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and further empirically learnable – MoL consistently outperforms dot products in HR across 6 scenarios by an average of 24.2% in HR@1 and 14.3% in HR@10, setting new state-of-the-art on recommendation datasets. We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds. Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. Given MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items [42, 3], our work concludes both theoretical and practical justifications for migrating vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.